A particle inside the rough surface of a rotating cone about its axis is at rest, relative to it as a height of 1m above its vertex. Friction coefficient is μ=0.5. If half angle of cone is 45°, the maximum angular velocity of revolution of cone can be :
1
Expert's answer
2020-08-02T15:10:46-0400

As per the question,

Height "(h)=1m"

Friction coefficient "(\mu) = 0.5"

Angle of cone "(\theta ) =45^\circ"

Let the maximum angular velocity is "\omega"

now,


Hence, radius "(r)" of the rotating circle,

"\tan(45^\circ)=\frac{r}{h}"

"r=h\tan (45^\circ) = 1"

So, "N= mg\cos 45^\circ+ mr\omega^2 \cos 45^\circ"

"=\frac{mg}{\sqrt{2}}+\frac{mr\omega^2}{\sqrt{2}}"

"=\frac{mg+mr\omega^2}{\sqrt{2}}"

"=\frac{m(g+r\omega^2)}{\sqrt{2}}"

So, "f_s =\mu N = \frac{0.5\times m(g+r\omega^2)}{\sqrt{2}}=\frac{m(g+r\omega^2)}{2\sqrt{2}}"

Hence, "mg\sin 45^\circ =\frac{m(g+r\omega^2)}{2\sqrt{2}}"

"\Rightarrow 2g=g+ r\omega^2"

"\Rightarrow \omega^2=g"

"\Rightarrow \omega =\sqrt{10}=\sqrt{9.8} rad/sec"